OpenStudy (ezraweise):

What is the graph of x + 3y = –9?

OpenStudy (ezraweise):

Thanks for Coming guys…

OpenStudy (ezraweise):

be right back

OpenStudy (ezraweise):

back

OpenStudy (ezraweise):

ok

OpenStudy (ezraweise):

Which is the equation of the line? A. 2x + 8y = 10 B. 2x – 8y = 8 C. 2x + 8y = 16 D. 2x – 8y = 16

OpenStudy (ezraweise):

?

OpenStudy (eli_moses):

The graph shows\[y=-\frac{ 1 }{ 4 }x+2\]

OpenStudy (eli_moses):

so...

OpenStudy (ezraweise):

what

OpenStudy (eli_moses):

im down to C and D

OpenStudy (ezraweise):

k

OpenStudy (eli_moses):

Its C

OpenStudy (ezraweise):

ok